# Exploring Holomorphic Retracts

### v6

@@ Conclusion
We now get to showing the existence of retracts of all possible dimensions through each point in this domain. To this end, pick any $c=(c_{1},c_{2},\ldots,c_{N})\in H^{N}$ and any positive integer $k$ less than $N$; we shall show that there exists a $k$-dimensional retract of $H^{N}$ passing through $c$. Consider the affine subspace $A=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:z_{j}=c_{j}~{}\text{for}~{}k+1\leq j\leq N\}$, defined by setting an appropriate number of components equal to that of the fixed point $c$. It suffices to prove that the $k$-dimensional affine subspace $H^{N}\cap A$ is a retract of $H$. First we do this for $k=N-1$ and the rest follows recursively as noted below. To this end, we let $L_{c_{N}}:=\{(z_{1},\ldots,z_{N-1},z_{N})\in\mathbb{C}^{N}:z_{N}=c_{N}\}$ and show that the $(N-1)$-dimensional affine subspace $H^{N}\cap L_{c_{N}}$ of $H^{N}$ is a retract of $H^{N}$. We split into $2$ cases depending on $c_{N}$ being zero or not. If $c_{N}=0$, then the linear projection map $\pi(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{N-1},0)$ suffices to render ${\rm image}(\pi)=H^{N}\cap L_{c_{N}}$ as a retract of $H^{N}$. If $c_{N}\neq 0$, then we consider the holomorphic map $\rho$ on $H^{N}$ defined by
@@
At this juncture, we can settle a question about whether holomorphic retraction mappings are automatically submersions so as (for instance) to render the hypotheses of theorem 1.17 redundant; to answer this, it suffices to consider the two-dimensional version of the above domain.
@@
Let $\rho=(\rho_{1},\rho_{2})$ be any of the retractions which map $H^{2}$ onto $Z$. Since second component of any point in $Z$ is $1$, we have $\rho_{2}\equiv 1$. No such drastic conclusion cannot be made for the first component of $\rho$ namely $\rho_{1}$, as we only have to begin with that it is a holomorphic function which maps $H^{2}$ onto $\Delta$; nevertheless we may deduce about its form as a consequence of certain features of $H^{2}$. To do so, consider first its restriction: $f(w):=\rho_{1}(0,w)$. Being an entire function which maps into $\Delta$, Liouville’s theorem implies that $f$ is a constant function. As $f(1)=0$, that constant is zero, thereby $f(w)=\rho_{1}(0,w)=0$ for all $w\in\mathbb{C}$. As $H^{2}$ is a logarithmically convex complete Reinhardt domain, $\rho_{1}$ can be expressed as a single power series expansion on $H^{2}$. This means that there exist scalars $c_{j,k}$ such that for all $(z,w)\in H^{2}$,
@@
Since $\rho_{1}(0,w)=0$ for all $w\in\mathbb{C}$, we obtain that $P_{2}\equiv 0$. Similarly we can consider another entire function $g:\mathbb{C}\to\Delta$ defined by $g(z)=\rho_{1}(z,0)$. By the same arguments, we obtain that $g$ for all $z\in\mathbb{C}$ and hence $P_{1}\equiv 0$. Therefore from the above equation (5.8), we arrive at our aforementioned deduction about the form of the first component namely, $\rho_{1}(z,w)=zwh(z,w)$ for some holomorphic function $h$ on $H^{2}$. We may now prove that $\rho$ is not submersive at origin by directly writing down the expression for its Jacobian matrix, which at any point $(z,w)$ is if the form
@@
A direct consequence of the above theorem is that unless we are in very special settings even within the class of balanced pseudoconvex domains, the hypothesis about submersivity of retractions as in theorem 1.17 cannot be dropped. For a quick (counter-)example, observe that the retract $Z$ in the foregoing theorem 5.4 cannot be realized as the graph over any linear retract of $H^{2}$ as in the conclusion of theorem 1.17. Indeed, the only (non-trivial) linear retracts of $H^{2}$ are the pair of coordinate axes, on neither of which $Z$ can be expressed as a graph, for then $Z$ would be biholomorphic to $\mathbb{C}$, manifestly false. It is also possible to give a non-affine example in this simple domain itself to show that there are retracts which are impossible to be realized as a graph over any affine linear subspace of $H$, thereby showing (again) that the conclusion of theorem 1.17 can fail as soon as its hypotheses, particularly about submersivity of the retractions fail which is the purpose of the following corollary.
@@
Suppose to arrive at a contradiction that $\tilde{\rho}:H^{2}\to H^{2}$ is a retraction map with ${\rm image}(\tilde{\rho})=\tilde{Z}$ and such that it is a submersion at the origin. Consider the automorphism $\psi:H^{2}\to H^{2}$ defined by $\psi(z,w)=(ze^{-zw},we^{zw})$. Conjugating $\rho$ by $\psi$, we then have the holomorphic map $\rho:H^{2}\to H^{2}$ given by $\rho(z,w)=\psi^{-1}\circ\tilde{\rho}\circ\psi(z,w)$. Note that $\rho$ is a retraction map on $H^{2}$ and ${\rm image}(\rho)=Z=\{(\zeta,1):\zeta\in\Delta\}$ and we may apply theorem 5.4 to $\rho$. But then the submersitivity of $\tilde{\rho}$ at the origin is equivalent to that of $\rho$ at the origin, as $\psi$ is not just a conjugating automorphism but also fixes at the origin; this contradicts theorem 5.4 finishing the proof. ∎
We next record a pair of bounded modifications of $H^{2}$ in the above examples wherein the some of the results hold as is, but some arguments cease to.
@@
In this case, by the aforementioned decomposition of $\partial\Omega$, we first note that the boundary point $(a^{\prime},b)$ lies in $\partial\Delta^{2}\cap\overline{H^{2}_{1/4}}$. The map $\rho$ on $\Delta^{2}$ defined by $\rho(z,w)=(z,b)$ suffices to give a holomorphic retraction map on $\Delta^{2}$ with ${\rm image}(\rho)=Z$. However, we must check that $Z$ is a retract of $\Omega$, not just of $\Delta^{2}$. To do this of-course we need only consider the restriction $\rho|_{\Omega}$, which maps $\Omega$ into $Z$ and that would finish the proof provided we verify the following.
@@
It now follows immediately that $\rho$ maps $\Omega$ into itself and ${\rm image}(\rho|_{\Omega})\subset L\cap\Omega=Z$; indeed, as we have for every $z\in Z=L\cap\Omega$ that $\rho(z)=z$, we have ${\rm image}(\rho|_{\Omega})=L\cap\Omega$. This finishes the proof completely that $L\cap\Omega$ is a retract of $\Omega$ passing through $(a,b)$ in the above case.
@@
In this case, again by the aforementioned decomposition of $\partial\Omega$, we start by noting that the boundary point $(a^{\prime},b)$ lies in $\overline{\Delta^{2}}\cap\partial H^{2}_{1/4}$. Note that $b\neq 0$ as $(a^{\prime},b)\in\partial H^{2}_{1/4}$. Now in contrast to the previous case, no linear choices work (the proof of which is easy but we shall not digress into that here), thereby leading us to consider the non-linear map $\varphi:H^{2}_{1/4}\to H^{2}_{1/4}$ defined by $\varphi(z,w)=(\frac{1}{b}zw,b)$. which is of-course a holomorphic retraction map on $H^{2}_{1/4}$ with ${\rm image}(\varphi)=W$. It remains to realize $W$ as a retraction of our smaller bounded $\Omega$, rather than the unbounded $H^{2}_{1/4}$. Again we need only consider the restriction $\varphi|_{\Omega}$ (which maps $\Omega$ into $W$), provided we verify the following.
@@
As another difference from the unbounded $H^{2}$, thereby a purpose of the above $\Omega$, is that the sets: $S_{1}$ of unit vectors representing directions along which there are linear retracts in $\Omega$, as well as, $S_{2}$ of those unit vectors along which there are no non-trivial retracts through the origin, are both non-trivial subsets of the unit sphere neither of which have measure zero (and with both having non-empty interior in the sphere).
@@
Example: Consider the domain $\Omega:=\mathbb{B}\cap H^{2}_{1/4}$, where $\mathbb{B}:=\{(z,w)\in\mathbb{C}^{2}:|z|^{2}+|w|^{2}<1\}$ be the euclidean ball centred at 00 of radius $1$ and $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}:|zw|<1/4\}$
@@
Let $L$ denotes the hyperplane $\{(z,w)\in\mathbb{C}^{2}:w=b\}$ Firstly note that
@@
Note that $\partial(L\cap\Omega)=L\cap\partial\Omega$ First we prove the following claim:
@@
Proof of the claim-1: If possible assume that $\partial(L\cap\Omega)\nsubseteq\partial\mathbb{B}$ and $\partial(L\cap\Omega)\nsubseteq\partial H^{2}_{1/4}$ Then there exist points $(a_{1},b)$ and $(b_{1},b)$ in $\partial(L\cap\Omega)$ such that $(a_{1},b)\notin\partial\mathbb{B}$ and $(b_{1},b)\notin\partial H^{2}_{1/4}$. From the above decomposition of $\partial\Omega$ and $(a_{1},b),(b_{1},b)\in\partial\Omega$, we obtain that $(a_{1},b)\in\mathbb{B}\cap\partial H^{2}_{1/4}$ and $(b_{1},b)\in\partial\mathbb{B}\cap H^{2}_{1/4}$. This implies that $a_{1},b_{1},b$ satisfies the following equations:
@@
Dividing $|a_{1}b|$ (which is equal to $1/4$) on both sides of the equation $|b_{1}b|<1/4$, we get $|b_{1}|<|a_{1}|$. Squaring and adding $|b|^{2}$ on both sides of the equation $|b_{1}|<|a_{1}|$, we get that $11=|b_{1}|^{2}+|b|^{2}<|a_{1}|^{2}+|b|^{2}<$ 1, which is vividly a contradiction. This finishes the proof of the claim.
@@
Since $(a^{\prime},b)\in\partial\Omega$ and the above decomposition of $\partial\Omega$, we obtain that $(a^{\prime},b)\in\partial\mathbb{B}\cap\overline{H^{2}_{1/4}}$. We know that $Z$ is an retract of $\mathbb{B}$. Then there exist a holomorphic retraction map $\rho$ on $\mathbb{B}$ such that $\rho(\mathbb{B})=Z$. Now we prove that $Z$ is a retract of $\Omega$. To prove this, consider the map $\rho|_{\Omega}$, which maps $\Omega$ into $Z$. Next we prove the following claim:
@@
Using the above claim-2, $\rho$ maps $\Omega$ into itself and $image(\rho|_{\Omega})\subset L\cap\Omega=Z$. If $z\in Z=L\cap\Omega$, then $\rho(z)=z$. So $image(\rho|_{\Omega})=L\cap\Omega$. Hence $L\cap\Omega$ is a retract of $\Omega$ passing through $(a,b)$.
@@
Note that $\varphi$ is a retraction map on $H^{2}_{1/4}$. Now we consider the map $\varphi|_{\Omega}$, which maps $\Omega$ into $W$. Next we prove the following claim:
@@
Proof of claim-3: First we prove that $|a^{\prime}|\geq\frac{1}{4|b|}$. If $|a^{\prime}|<\frac{1}{4|b|}$ then $|a^{\prime}b|<1/4$ and thus implies that $(a^{\prime},b)\in H^{2}_{1/4}$, which contradicts the fact that $(a^{\prime},b)\in\partial H^{2}_{1/4}$ Hence we proved that $|a^{\prime}|\geq\frac{1}{4|b|}$
@@
Let us firstly quickly note that if $D$ is convex near $p$ is the sense of definition 2.30, then $D$ is geometrically convex near $p$ i.e., there exist an open ball $U$ in $\mathbb{C}^{N}$ centered at $p$ such that $U\cap D$ is convex in the usual sense of being closed under the formation of line segments. Indeed, this follows by considering the closed half-spaces (containing the domain) given by the linear functionals as per definition 2.30 due to the points of $\partial D\cap U$ being convex boundary points; supporting hyperplanes (and likewise half-spaces containing $D\cap U$) at other points of $\partial(D\cap U)$ exist just by convexity of the ball $U$. As the intersection of these closed half-spaces precisely equals $\overline{U\cap D}$, the convexity of $U\cap D$ follows.
Next, we show the subadditivity of $h$. Let $z,w\in U\cap D$. By definition of the Minkowski functional, there exist positive reals $t,s$ such that $z/t,w/s\in U\cap D$ and $t<h(z)+\epsilon$, $s<h(w)+\epsilon$ for some $\epsilon>0$. Writing $(z+w)/(s+t)$ as the convex combination:
$\frac{z+w}{s+t}=\frac{t}{s+t}\cdot\frac{z}{t}+\frac{s}{s+t}\cdot\frac{w}{s},$
it follows that by the convexity of $U\cap D$, that $(z+w)/(s+t)\in U\cap D$. Hence $h(z+w)\leq s+t<h(z)+h(w)+2\epsilon$. As $\epsilon$ is arbitrary, we get $h(z+w)\leq h(z)+h(w)$ for all $z,w\in U\cap D$, finishing the verification of the required subadditivity. We now proceed to the original goal about continuity of $h$. Let $q$ be arbitrary point in $U\cap\partial D$ and $(z_{n})$ be a sequence converging to $q\in U\cap\partial D$. (without loss of generality, we assume that $z_{n}\in U\cap D$ for all $n$). By subadditivity of $h$ and the well-known basic fact that the Minkowski functional is dominated by some scaling of the Euclidean norm, we have $|h(z_{n})-h(q)|\leq h(z_{n}-q)\leq c\|z_{n}-q\|$ for some $c>0$. The convergence of $z_{n}$ to $q$, thereby readily implies that $h(z_{n})$ converges to $h(q)$. Hence $h$ is continuous near $p$, finishing the proof owing to what was noted earlier. ∎
@@
Consider the holomorphic map $\varphi:\Delta\to\mathbb{C}^{N}$ defined by $\varphi(t)=(p+tu)t$. As $p+tu\in\partial D$ for all $t\in\Delta$ and $h$ is continuous at $p+tu$ for all $t\in\Delta$, we have by a well-known fundamental fact (see for instance Remark 2.2.1 (h) of [Jrncki_invrnt_dst]) that $h(p+tu)=1$ for all $t\in\Delta$. Hence
@@
So, $\varphi$ maps $\Delta$ into $D$. Next we want to prove that $\varphi$ is a complex geodesic in $D$. First we claim that there exist $t_{0}\in\Delta^{*}:=\Delta\setminus\{0\}$ such that $D$ is convex at $p+t_{0}u$. By the definition of convexity of $\partial D$ at $p$, there exist a $\mathbb{C}$-linear functional $L$ such that
@@
Since $\mathbb{C}^{N}=Ker(L)\oplus span\{p\}$, we can write $u=u_{1}+u_{2}$, where $u_{1}\in Ker(L)$ and $u_{2}\in span\{p\}$. Write $u_{2}=\lambda p$ for some $\lambda\in\mathbb{C}$. So
@@
We need only prove that there exist $t_{0}\in\Delta^{*}$ such that $|1+\lambda t_{0}|=1$, which we do now. If $\lambda=0$, then $u\in Ker(L)$. Hence (6.1) holds for any $t_{0}\in\Delta^{*}$. Assume that $\lambda\neq 0$. Write $\lambda t=re^{i\theta}$ where $\theta\in[-\pi,\pi)$. From this we get $t=\frac{1}{\lambda}(re^{i\theta})$. The relations $|1+\lambda t|=1$ and $|t|<1$ becomes $|1+re^{i\theta}|=1$ and $|re^{i\theta}|<|\lambda|$. We want to prove that there exist $(r,\theta)$ such that
@@
Geometrically above equation represents the intersection of circle centred at $-1$ with radius $1$ and the disc centred at 00 of radius $|\lambda|$. So there exist $t_{0}\in\Delta^{*}$ such that $|1+\lambda t_{0}|=1$. This proves the claim.
@@
To proceed further with the proof that $\varphi$ is a complex $C$-geodesic, let us now note that as $D$ is convex at the boundary point $p+t_{0}u$ by the foregoing, we have by proposition 2.3.1(b) in [Jrncki_invrnt_dst], with the same notations therein, that
@@
By the hypotheses that $D$ is bounded and from remark 2.2.1 (l) in [Jrncki_invrnt_dst], we obtain that $h(u)\neq 0$. Hence we get $|\lambda|=1/h(u)$. We now split into two cases depend upon whether $h(u)>1$ or not.

### v7

@@ Conclusion
We now get to showing the existence of retracts of all possible dimensions through each point in this domain. To this end, pick any $c=(c_{1},c_{2},\ldots,c_{N})\in H^{N}$ and any positive integer $k$ less than $N$; we shall show that there exists a $k$-dimensional retract of $H^{N}$ passing through $c$. Consider the affine subspace $A=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}:z_{j}=c_{j}~\text{for}~k+1\leq j\leq N\}$, defined by setting an appropriate number of components equal to that of the fixed point $c$. It suffices to prove that the $k$-dimensional affine subspace $H^{N}\cap A$ is a retract of $H$. First, we do this for k=N1k=N-1 and the rest follows recursively as noted below. To this end, we let $L_{c_{N}}:=\{(z_{1},\ldots,z_{N-1},z_{N})\in\mathbb{C}^{N}:z_{N}=c_{N}\}$ and show that the (N1)(N-1)-dimensional affine subspace $H^{N}\cap L_{c_{N}}$ of $H^{N}$ is a retract of $H^{N}$. We split into $2$ cases depending on $c_{N}$ being zero or not. If $c_{N}=0$, then the linear projection map $\pi(z_{1},\ldots,z_{N})=(z_{1},\ldots,z_{N-1},0)$ suffices to render ${\rm image}(\pi)=H^{N}\cap L_{c_{N}}$ as a retract of $H^{N}$. If $c_{N}\neq 0$, then we consider the holomorphic map $\rho$ on $H^{N}$ defined by
@@
At this juncture, we can settle a question about whether holomorphic retraction mappings are automatically submersions so as (for instance) towards rendering hypotheses such as (ii) in theorem 1.19 redundant in general. We briefly address this by considering the two-dimensional version of the above domain (we note here that $D_{L}$ is not bounded as stipulated by theorem 1.19).
@@
Let $\rho=(\rho_{1},\rho_{2})$ be any of the retractions which map $H^{2}$ onto $Z$. Since second component of any point in $Z$ is $1$, we have $\rho_{2}\equiv 1$. No such drastic conclusion cannot be made for the first component of $\rho$ namely $\rho_{1}$, as we only have to begin with that it is a holomorphic function which maps $H^{2}$ onto $\Delta$; nevertheless we may deduce about its form as a consequence of certain features of $H^{2}$. To do so, consider first its restriction: $f(w):=\rho_{1}(0,w)$. Being an entire function which maps into $\Delta$, Liouville’s theorem implies that $f$ is a constant function. As $f(1)=0$, that constant is zero, thereby $f(w)=\rho_{1}(0,w)=0$ for all $w\in\mathbb{C}$. As $H^{2}$ is a (logarithmically convex) complete Reinhardt domain, $\rho_{1}$ can be expressed as a single power series expansion on $H^{2}$. This means that there exist scalars $c_{j,k}$ such that for all $(z,w)\in H^{2}$,
@@
Since $\rho_{1}(0,w)=0$ for all $w\in\mathbb{C}$, we obtain that $P_{2}\equiv 0$. Similarly we can consider another entire function $g:\mathbb{C}\to\Delta$ defined by $g(z)=\rho_{1}(z,0)$. By the same arguments, we obtain that $P_{1}\equiv 0$. Therefore from the above equation (5.10), we arrive at our aforementioned deduction about the form of the first component namely, $\rho_{1}(z,w)=zwh(z,w)$ for some holomorphic function $h$ on $H^{2}$. We may now prove that $\rho$ is not submersive at origin by directly writing down the expression for its Jacobian matrix, which at any point $(z,w)$ is if the form
@@
A direct consequence of the above theorem is that unless we are in very special settings even within the class of balanced pseudoconvex domains, the hypothesis (ii) along-with boundedness of $D_{L}$ as in theorem 1.19 cannot be dropped. For a quick (counter-)example, observe that the retract $Z$ in the foregoing theorem 5.13 cannot be realized as the graph over any linear retract of $H^{2}$ as in the conclusion of theorem 1.19. Indeed, the only (non-trivial) linear retracts of $H^{2}$ are the pair of coordinate axes, on neither of which $Z$ can be expressed as a graph, for then $Z$ would be biholomorphic to $\mathbb{C}$, manifestly false. It is also possible to give a non-affine example in this simple domain itself to show that there are retracts which are impossible to be realized as a graph over any affine linear subspace of $H$, thereby showing (again) that the conclusion of theorem 1.19 can fail as soon as some of its hypotheses is dropped as mentioned above, which is the purpose of the following corollary.
@@
Suppose to arrive at a contradiction that $\tilde{\rho}:H^{2}\to H^{2}$ is a retraction map with ${\rm image}(\tilde{\rho})=\tilde{Z}$ and such that it is a submersion at the origin. Consider the automorphism $\psi:H^{2}\to H^{2}$ defined by $\psi(z,w)=(ze^{-zw},we^{zw})$. Conjugating $\rho$ by $\psi$, we then have the holomorphic map $\rho:H^{2}\to H^{2}$ given by $\rho(z,w)=\psi^{-1}\circ\tilde{\rho}\circ\psi(z,w)$. Note that $\rho$ is a retraction map on $H^{2}$ and ${\rm image}(\rho)=Z=\{(\zeta,1):\zeta\in\Delta\}$ and we may apply theorem 5.13 to $\rho$. But then the submersivity of $\tilde{\rho}$ at the origin is equivalent to that of $\rho$ at the origin, as $\psi$ is not just a conjugating automorphism but also fixes at the origin; this contradicts theorem 5.13, finishing the proof. ∎
We next record a pair of bounded modifications of $H^{2}$ in the above examples wherein the some of the results hold as is, but some arguments cease to. More importantly, these give examples of bounded balanced non-convex (but pseudoconvex) domains wherein (i) for each point, there is a (non-trivial) retract passing through it and (ii) they are equivalent to product domains but contain an open piece of directions (an open subset of the unit sphere $\partial\mathbb{B}$ representing directions) along which retract exist.
@@
In this case, by the aforementioned decomposition of $\partial\Omega$, we first note that the boundary point $(a^{\prime},b)$ lies in $\partial\Delta^{2}\cap\overline{H^{2}_{1/4}}$. The map $\rho$ on $\Delta^{2}$ defined by $\rho(z,w)=(z,b)$ suffices to give a holomorphic retraction map on $\Delta^{2}$ with ${\rm image}(\rho)=Z$. However, we must check that $Z$ is a retract of $\Omega$, not just of $\Delta^{2}$. To do this of-course we need only consider the restriction $\rho_{|_{\Omega}}$, which maps $\Omega$ into $Z$ and that would finish the proof provided we verify the following.
@@
It now follows immediately that $\rho$ maps $\Omega$ into itself and ${\rm image}(\rho_{|_{\Omega}})\subset L\cap\Omega=Z$; indeed, as we have for every $z\in Z=L\cap\Omega$ that $\rho(z)=z$, we have ${\rm image}(\rho_{|_{\Omega}})=L\cap\Omega$. This finishes the proof completely that $L\cap\Omega$ is a retract of $\Omega$ passing through $(a,b)$ in the above case.
@@
In this case, again by the aforementioned decomposition of $\partial\Omega$, we start by noting that the boundary point $(a^{\prime},b)$ lies in $\overline{\Delta^{2}}\cap\partial H^{2}_{1/4}$. Note that $b\neq 0$ as $(a^{\prime},b)\in\partial H^{2}_{1/4}$. Now in contrast to the previous case, no linear choices work (the proof of which is easy but we shall not digress into that here), thereby leading us to consider the non-linear map $\varphi:H^{2}_{1/4}\to H^{2}_{1/4}$ defined by $\varphi(z,w)=(\frac{1}{b}zw,b)$. which is of-course a holomorphic retraction map on $H^{2}_{1/4}$ with ${\rm image}(\varphi)=W$. It remains to realize $W$ as a retraction of our smaller bounded $\Omega$, rather than the unbounded $H^{2}_{1/4}$. Again we need only consider the restriction $\varphi_{|_{\Omega}}$ (which maps $\Omega$ into $W$), provided we verify the following.
@@
As another difference from the unbounded $H^{2}$, thereby a purpose of the above $\Omega$, is that the sets: $S_{1}$ of unit vectors representing directions along which there are linear retracts in $\Omega$, as well as, $S_{2}$ of those unit vectors along which there are no non-trivial retracts through the origin, are both non-trivial subsets of the unit sphere neither of which have measure zero (and with both having non-empty interior in the sphere). Finally here, we also mention that there are non-linear retracts along the directions given by $S_{1}$ through the origin. Thus there exist bounded balanced pseudoconvex domains which are not biholomorphic to product domains but still have non-linear retracts through the origin. Indeed, we may establish $D:=\Delta^{2}\cap H^{1/2}$ is irreducible as follows. Assuming the contrary, we observe that it would follow by the Riemann mapping theorem that $D$ is biholomorphic to the bidisc $\Delta^{2}$. The homogeneity of $\Delta^{2}$ together with an application of Cartan’s theorem then renders a linear equivalence between $\Delta^{2}$ and $D$ which is impossible as the former is convex whereas the latter is not.
@@
Example: Consider the domain $\Omega:=\mathbb{B}\cap H^{2}_{1/4}$, where $\mathbb{B}:=\{(z,w)\in\mathbb{C}^{2}:|z|^{2}+|w|^{2}<1\}$ be the euclidean ball centred at 0 of radius $1$ and $H^{2}_{1/4}=\{(z,w)\in\mathbb{C}^{2}:|zw|<1/4\}$ .
@@
Let $L$ denotes the hyperplane $\{(z,w)\in\mathbb{C}^{2}:w=b\}$ . Firstly note that
@@
Note that $\partial(L\cap\Omega)=L\cap\partial\Omega$ . First we prove the following claim:
@@
Proof of the claim-1: If possible assume that $\partial(L\cap\Omega)\nsubseteq\partial\mathbb{B}$ and $\partial(L\cap\Omega)\nsubseteq\partial H^{2}_{1/4}$ . Then there exist points $(a_{1},b)$ and $(b_{1},b)$ in $\partial(L\cap\Omega)$ such that $(a_{1},b)\notin\partial\mathbb{B}$ and $(b_{1},b)\notin\partial H^{2}_{1/4}$. From the above decomposition of $\partial\Omega$ and $(a_{1},b),(b_{1},b)\in\partial\Omega$, we obtain that $(a_{1},b)\in\mathbb{B}\cap\partial H^{2}_{1/4}$ and $(b_{1},b)\in\partial\mathbb{B}\cap H^{2}_{1/4}$. This implies that $a_{1},b_{1},b$ satisfies the following equations:
@@
Dividing $|a_{1}b|$ (which is equal to $1/4$) on both sides of the equation $|b_{1}b|<1/4$, we get $|b_{1}|<|a_{1}|$. Squaring and adding $|b|^{2}$ on both sides of the equation $|b_{1}|<|a_{1}|$, we get that $1=|b_{1}|^{2}+|b|^{2}<|a_{1}|^{2}+|b|^{2}<1$, which is vividly a contradiction. This finishes the proof of the claim.
@@
Since $(a^{\prime},b)\in\partial\Omega$ and the above decomposition of $\partial\Omega$, we obtain that $(a^{\prime},b)\in\partial\mathbb{B}\cap\overline{H^{2}_{1/4}}$. We know that $Z$ is an retract of $\mathbb{B}$. Then there exist a holomorphic retraction map $\rho$ on $\mathbb{B}$ such that $\rho(\mathbb{B})=Z$. Now we prove that $Z$ is a retract of $\Omega$. To prove this, consider the map $\rho_{|_{\Omega}}$, which maps $\Omega$ into $Z$. Next we prove the following claim:
@@
Using the above claim-2, $\rho$ maps $\Omega$ into itself and $image(\rho_{|_{\Omega}})\subset L\cap\Omega=Z$. If $z\in Z=L\cap\Omega$, then $\rho(z)=z$. So $image(\rho_{|_{\Omega}})=L\cap\Omega$. Hence $L\cap\Omega$ is a retract of $\Omega$ passing through $(a,b)$.
@@
Note that $\varphi$ is a retraction map on $H^{2}_{1/4}$. Now we consider the map $\varphi_{|_{\Omega}}$, which maps $\Omega$ into $W$. Next we prove the following claim:
@@
Proof of claim-3: First we prove that $|a^{\prime}|\geq\frac{1}{4|b|}$. If $|a^{\prime}|<\frac{1}{4|b|}$ then $|a^{\prime}b|<1/4$ and thus implies that $(a^{\prime},b)\in H^{2}_{1/4}$, which contradicts the fact that $(a^{\prime},b)\in\partial H^{2}_{1/4}$ . Hence we proved that $|a^{\prime}|\geq\frac{1}{4|b|}$
@@
Let us firstly quickly note that if $D$ is convex near $p$ in the sense of definition 2.34, then $D$ is geometrically convex near $p$ i.e., there exist an open ball $U$ in $\mathbb{C}^{N}$ centered at $p$ such that $U\cap D$ is convex in the usual sense of being closed under the formation of line segments. Indeed, this follows by considering the closed half-spaces (containing the domain) given by the linear functionals as per definition 2.34 due to the points of $\partial D\cap U$ being convex boundary points; supporting hyperplanes (and likewise half-spaces containing $D\cap U$) at other points of $\partial(D\cap U)$ exist just by convexity of the ball $U$. As the intersection of these closed half-spaces precisely equals $\overline{U\cap D}$, the convexity of $U\cap D$ follows.
Let $C:=\bigcup_{z\in\partial D\cap U}\{\lambda z:\lambda\geq 0\}$ and $C_{D}:=C\cap D$. As $U\cap D$ is geometrically convex, so is the cone generated by it, thereby assuring that $C_{D}$ is convex. Next, we show the subadditivity of $h$ on $C$. Let $z,w\in C_{D}$. By definition of the Minkowski functional, for each $\epsilon>0$ there exist positive reals $t,s$ such that $z/t,w/s\in D$ and $t<h(z)+\epsilon$, $s<h(w)+\epsilon$. Since $C_{D}$ is open, there exist $t^{\prime},s^{\prime}$ such that $t^{\prime}<t,~s^{\prime}<s$ and $z/t^{\prime},w/s^{\prime}\in C_{D}$. Writing $(z+w)/(s^{\prime}+t^{\prime})$ as the convex combination:
$\frac{z+w}{s^{\prime}+t^{\prime}}=\frac{t^{\prime}}{s^{\prime}+t^{\prime}}\cdot\frac{z}{t^{\prime}}+\frac{s^{\prime}}{s^{\prime}+t^{\prime}}\cdot\frac{w}{s^{\prime}},$
it follows that by the convexity of $C_{D}$, that $(z+w)/(s+t)\in C_{D}$. Hence $h(z+w)\leq s^{\prime}+t^{\prime}<h(z)+h(w)+2\epsilon$. As $\epsilon$ is arbitrary, we get $h(z+w)\leq h(z)+h(w)$ for all $z,w\in C_{D}$. This shows that $h$ is sub-additive on $C_{D}$. To prove the subadditivity of $h$ on $C$, let $z,w$ be any two points in $C$. Choose $k$ so that $k>\max\{h(z),h(w)\}$. Note that $z/k,w/k\in C_{D}$. Applying the subadditivity of $h$ on $C_{D}$ and homogenity of $h$, we obtain that $h(z/k+w/k)\leq 1/k~h(z)+1/k~h(w)$. This shows that $h$ is subadditive on $C$. The subadditivity of $h$ on $C$ and together with homogenity of $h$ yields that $h$ is a convex function on $C$. Hence $h$ is continuous on $C$ and in particular $h$ is continuous near $p$. ∎
@@
Consider the holomorphic map $\varphi:\Delta\to\mathbb{C}^{N}$ defined by $\varphi(t)=(p+tu)t$. As $p+tu\in\partial D$ for all $t\in\Delta$ and $h$ is continuous at $p+tu$ for all $t\in\Delta$, we have by a well-known fundamental fact (see proposition 2.9 (d)) that $h(p+tu)=1$ for all $t\in\Delta$. Hence
@@
So, $\varphi$ maps $\Delta$ into $D$. Next we want to prove that $\varphi$ is a complex $C$-geodesic in $D$. First we claim that there exist $t_{0}\in\Delta^{*}:=\Delta\setminus\{0\}$ such that $D$ is convex at $p+t_{0}u$. By the definition of convexity of $\partial D$ at $p$, there exist a $\mathbb{C}$-linear functional $L$ such that
@@
Since $\mathbb{C}^{N}=\ker(L)\oplus span\{p\}$, we can write $u=u_{1}+u_{2}$, where $u_{1}\in\ker(L)$ and $u_{2}\in span\{p\}$. Write $u_{2}=\lambda p$ for some $\lambda\in\mathbb{C}$. So
@@
We need only prove that there exist $t_{0}\in\Delta^{*}$ such that $|1+\lambda t_{0}|=1$, which we do now. If $\lambda=0$, then $u\in\ker(L)$. Hence (6.1) holds for any $t_{0}\in\Delta^{*}$. Assume that $\lambda\neq 0$. Write $\lambda t=re^{i\theta}$ where $\theta\in[-\pi,\pi)$. From this we get $t=\frac{1}{\lambda}(re^{i\theta})$. The relations $|1+\lambda t|=1$ and $|t|<1$ becomes $|1+re^{i\theta}|=1$ and $|re^{i\theta}|<|\lambda|$. We want to prove that there exist $(r,\theta)$ such that
@@
Geometrically above equation represents the intersection of circle centred at 1-1 with radius $1$ and the disc centred at 0 of radius $|\lambda|$. So there exist $t_{0}\in\Delta^{*}$ such that $|1+\lambda t_{0}|=1$. This proves the claim.
@@
To proceed further with the proof that $\varphi$ is a complex $C$-geodesic, let us now note that as $D$ is convex at the boundary point $p+t_{0}u$ by the foregoing; thereby also at $(t_{0}/|t_{0}|)(p+t_{0}u)$. We then have by proposition 2.3.1(b) in [Jrncki_invrnt_dst], with the same notations therein, that
@@
By the hypotheses that $D$ is bounded and from proposition 2.9 (f), we obtain that $h(u)\neq 0$. Hence we get $|\lambda|=1/h(u)$. We now split into two cases depend upon whether $h(u)>1$ or not.